OpenStudy (anonymous):

divide /simplify 2x-6/121 divide by x-3/55x

OpenStudy (anonymous):

\[\frac{2x-6}{121}\div \frac{x-3}{55x}?\]

OpenStudy (anonymous):

\[\frac{10x}{11}\]

OpenStudy (anonymous):

10x/11